# Discrete Duality Symmetry in a $3D$ Field-Theoretic Model

### v2

@@
Abstract: We demonstrate the discrete duality symmetry between the Abelian 1-form and 2-form basic gauge fields in the context of a three $(2+1)$-dimensional ($3D$) combined system of the field-theoretic model for the combined system of the free Abelian 1-from and 2-form gauge theories within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism where the classical gauge-fixed Lagrangian density of this theory is generalized to its quantum counterpart as the BRST and co-BRST invariant Lagrangian density. We show clearly the existence of the off-shell nilpotent (co-)BRST symmetry transformations and establish their intimate connection through a set of underlying discrete duality symmetry transformations in our BRST-quantized theory. We provide the mathematical basis for the existence of the discrete duality symmetry transformations in our theory through the Hodge duality operator (that is defined on the $3D$ flat Minkowskian spacetime manifold). We briefly mention a bosonic symmetry transformation which is constructed from the anticommutator of the above off-shell nilpotent (co-)BRST symmetry transformations. We lay emphasis on the algebraic structures of the existing continuous and discrete duality symmetry transformations in our theory (where they are treated as operators).
@@
Keywords: A $3D$ field-theoretic model; combined system of the Abelian 1-form and 2-form gauge theories; first-class constraints; (dual-)gauge symmetries; (anti-)BRST symmetries; (anti-)co-BRST symmetries; discrete duality symmetry transformations
@@ 1 Introduction
The study of the higher $p$-form ($p=2,3,4,...$) gauge theories***The most successful theory of THEP is the standard model of particle physics (SMPP) which is based on the interacting non-Abelian 1-form (i.e. $p=1$) gauge theory where there is a stunning degree of agreement between the theory and experiment. However, the conclusive experimental evidence of the masses of the neutrinos has compelled the modern-day practitioners of THEP to go beyond the domain of the validity of SMPP. The ideas behind the (super)string theories are very promising in this direction. has become important and essential due to the appearance of their basic fields in the quantum excitations of (super)strings which are the forefront areas of research activities in the realm of theoretical high energy physics (THEP). The modern developments in the domain of (super)string theories (see, e.g. [1-5] and references therein) have brought together, in an unprecedented manner, the active researchers in THEP and pure mathematics on a single intellectual platform where both kinds of researchers have benefited from the knowledge and creativity of each-other. In other words, there has been a notable convergence of ideas from the research activities in the realms of THEP and pure mathematics. Such kinds of confluence of ideas have been observed in our studies of the massless and Stückelberg-modified massive Abelian $p$-form (i.e. $p=1,2,3$) gauge theories (see, e.g. [6-9] and references therein) within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism [10-13] where it has been established that, in the $D=2p$ (i.e. $D=2,4,6$) dimensions of spacetime, these theories turn out to be the field-theoretic examples for Hodge theory. The discrete and continuous symmetry transformations (and corresponding conserved charges) of these theories have been able to provide the physical realizations of the de Rham cohomological operators of differential geometry (see, e.g. [14-18]) at the algebraic level. In other words, in our research activities, it has been established that there is a convergence of ideas from the physical aspects of the BRST formalism and the mathematical ingredients related to the cohomological operators of differential geometry. It is crystal clear, from our present discussions, that our studies [6-9] on the above BRST-quantized Abelian theories have been confined only to the even dimensions (i.e. $D=2,4,6$) of the flat Minkowskian spacetime.
@@
The existence of the discrete duality symmetry transformations in our $3D$ theory is very important as they provide (i) a beautiful algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformations (cf. Sec. 4), and (ii) a direct connection between the fermionic (i.e. nilpotent) (co-)BRST symmetry transformations of our $3D$ theory (cf. Appendix A). In our recent couple of papers [19,20], we have not been able to obtain such kinds of relationships due to the fact that we were unaware of the existence of such kinds of beautiful discrete duality symmetry transformations. As a consequence, we could not provide the physical realization of the Hodge duality operator (in our earlier works [19,20]) which plays a key role in the algebraic structures that are satisfied by the de Rham cohomological operators of differential geometry. In our present endeavor, we have given a great deal of importance to the discrete duality symmetry transformations because they provide (i) a deep connection between the Abelian 1-form and 2-form basic gauge fields [cf. Eq. (3)], and (ii) the physical realization of the Hodge duality operator [14-18] the relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives [i.e. $(\delta)d$].
@@
One of the key upshots of our present endeavor is the observation that there is existence of the massless (i.e. $\Box\widetilde{\phi}=0,\;\Box\phi=0$) (pseudo-)scalar fields [i.e. $(\widetilde{\phi})\phi$] in our theory which carry the kinetic terms that are endowed with the opposite signs. However, despite this obvious distinct difference, they obey the massless Klein-Gordon equations of motion: $\Box\widetilde{\phi}=0,\;\Box\phi=0$. It is interesting to point that the fields with negative kinetic terms have become quite popular in the realm of the cosmological models of the Universe. In our theory, on the symmetry grounds alone, we observe that the pseudo-scalar field ($\widetilde{\phi}$) carries the negative kinetic term. Such kinds of fields are a set of possible candidates for (i) the “phantom” and/or “ghost” fields in the realm of the cyclic, bouncing and self-accelerated cosmological models of the Universe [21-25], and (ii) the dark energy (see, e.g. [26,27] and references therein) because these fields automatically lead to the creation of the negative pressure in the theory (which happens to be one of the characteristic features of dark energy). On the contrary, the pure scalar field ($\phi$), with the positive kinetic term, is a normal field which does not lead to the existence of the negative pressure in the theory. Hence, the existence of the pseudo-scalar field (in our theory) is very interesting in the current scenario of research activities that are related to(i) the realms of the cosmological models of the Universe, and (ii) the search for the possible dark energy candidate(s).
@@
In our Appendix A, we establish a direct connection between the BRST and co-BRST symmetry transformations by exploiting the theoretical beauty and strength of the discrete duality symmetry transformations of our $3D$ theory.
@@ 2 Preliminaries: Constraint Analysis, (Dual-)Gauge and Discrete Duality Symmetry Transformations
We begin with the $3D$ field-theoretic model for the combined system of the free Abelian 1-form (i.e. $A^{(1)}=A_{\mu}\,dx^{\mu}$) and 2-form [i.e. $B^{(2)}=\frac{1}{2!}\,B_{\mu\nu}(dx^{\mu}\wedge dx^{\nu})$] gauge theories which is described by the following standard Lagrangian densityOur three $(2+1)$-dimensional ($3D$) Minkowskian spacetime is characterized by the flat metric tensor $\eta_{\mu\nu}=\text{diag}\,(+1,-1,-1)$ so that the dot product between two non-null Lorentz vectors $U_{\mu}$ and $V_{\mu}$ is defined as: $U\cdot V=\eta_{\mu\nu}\,U^{\mu}\,V^{\nu}\equiv U_{0}\,V_{0}-U_{i}\,V_{i}$ where the Greek indices $\mu,\nu,\sigma,...=0,1,2$ stand for the time and space directions and the Latin indices $i,j,k,...=1,2$ denote the space directions only. The $3D$ Levi-Civita tensor $\varepsilon_{\mu\nu\sigma}$ is chosen such that $\varepsilon_{012}=\varepsilon^{012}=+1$ and the standard relationships: $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\sigma}=+3!$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\eta}=+2!\,\delta^{\sigma}_{\eta}$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\eta\kappa}=+1!\,(\delta^{\nu}_{\eta}\,\delta^{\sigma}_{\kappa}-\delta^{\nu}_{\kappa}\,\delta^{\sigma}_{\eta})$ are satisfied. We also adopt the convention of the derivative w.r.t. the antisymmetric tensor field as: $\big{(}\partial B_{\mu\nu}/\partial B_{\sigma\kappa}\big{)}=\frac{1}{2!}\,(\delta^{\sigma}_{\mu}\,\delta^{\kappa}_{\nu}-\delta^{\sigma}_{\nu}\,\delta^{\kappa}_{\mu}),etc.$
@@
where the field-strength tensors: $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $H_{\mu\nu\sigma}=\partial_{\mu}B_{\nu\sigma}+\partial_{\nu}B_{\sigma\mu}+\partial_{\sigma}B_{\mu\nu}$ are derived from the 2-form $F^{(2)}=d\,A^{(1)}=\frac{1}{2!}\,F_{\mu\nu}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}$ and 3-form $H^{(3)}=d\,B^{(2)}=\frac{1}{3!}\,H_{\mu\nu\sigma}\,\big{(}dx^{\mu}\wedge dx^{\nu}\wedge dx^{\sigma}\big{)}$, respectively. Here the symbol $d=\partial_{\mu}\,dx^{\mu}$ [with $d^{2}=\frac{1}{2!}\,\big{(}\partial_{\mu}\,\partial_{\nu}-\partial_{\nu}\,\partial_{\mu}\big{)}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}=0$] denotes the exterior derivative of differential geometry (see, e.g. [14-18]) and the above 1-form (i.e. $A^{(1)}$) and 2-form (i.e. $B^{(2)}$) define the Lorentz vector gauge field $A_{\mu}$ and the second-rank antisymmetric (i.e. $B_{\mu\nu}=B_{\nu\mu}$) tensor gauge field $B_{\mu\nu}$, respectively. It is the specific feature of the three $(2+1)$-dimensional $(3D)$ Minkowskian spacetime that the kinetic terms for the Abelian 1-form gauge field $A_{\mu}$ and the antisymmetric tensor gauge field $B_{\mu\nu}$ can be expressed in terms of the $3D$ Levi-Civita tensor and the derivative on them as we have expressed these terms in equation (1). One of the key reasons behind this observation is the fact that, in the $3D$ Minkowskian spacetime, we have: $\frac{1}{12}\,H^{\mu\nu\sigma}\,H_{\mu\nu\sigma}=\frac{1}{2}\,H^{012}\,H_{012}\equiv\frac{1}{2}\,(H_{012})^{2}$. The component $H_{012}$, however, can be written in its covariant form as: $H_{012}=\frac{1}{2}\,\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}$ which has been taken into account in our equation (1) to express the kinetic term for the Abelian 2-form gauge field. Similarly, it is straightforward to check that the $3D$ kinetic term for the Abelian 1-form field is: $-\frac{1}{4}\,F^{\mu\nu}\,F_{\mu\nu}=-\frac{1}{2}\,\big{(}\varepsilon^{\mu\nu\sigma}\partial_{\nu}A_{\sigma}\big{)}^{2}$. For our combined system of the $3D$ free Abelian 1-form and 2-form gauge theories, the infinitesimal, local and continuous gauge transformations $(\delta_{g})$ are well-known. These symmetry transformations for our $3D$ system are:
@@
To determine these first-class constraints, the first-step is to derive the canonical conjugate momenta (i.e. $\Pi^{\mu\mu}_{(B)},\;\Pi^{\mu}_{(A)}$ ) w.r.t. the basic gauge fields $B_{\mu\nu}$ and $A_{\mu}$, respectively. For the $3D$ version of the Lagrangian density (1), these momenta are:
@@
The above equation demonstrates that we have two sets of primary constraints (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(B)}\approx 0$) on our theory where the symbol $\approx 0$ stands for weakly zero in Dirac’s notation (see, e.g. [28,29] for details). The requirements of the time-evolution invariance of these primary constraints lead to the determination of the secondary constraints. The following Euler-Lagrange (EL) equations of motion (EoMs)
@@
where $\Pi^{ij}_{(B)}=\frac{1}{2}\,\varepsilon^{0ij}\,H_{012}$ and $\Pi^{i}_{(A)}=-F^{0i}$ are the space components of the conjugate momenta that have been defined in equation (2). It is obvious that we have four sets of constraints on our theory out of which two sets (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(A)}\approx 0$) are the primary constraints and the rest of the two (i.e. $\partial_{j}\Pi^{ji}_{(B)}\approx 0,\;\partial_{i}\Pi^{i}_{(A)}\approx 0$) are the secondary constraints [28,29]. Since all these constraints are expressed in terms of the components of the conjugate momenta (2), all of them commute among themselves. Hence, they belong to the first-class variety according to Dirac’s prescription for the classification scheme of constraints [16,17]. These first-class constraints generate the infinitesimal gauge symmetry transformations (2) which we explain in the following paragraph.
According to Noether’s theorem, the existence of the continuous symmetry transformations always leads to the derivations of the Noether conserved currents and corresponding charges. Due to the infinitesimal gauge symmetry transformations in equation (2), we find the following expression for the Noether gauge current $J^{\mu}_{(G)}$ for our $3D$ field-theoretic combined system of the free Abelian 1-form and 2-form theory
@@
Since we have constraints (i.e. $-F^{00}\approx 0,\;\frac{1}{2}\,\varepsilon^{00i}\,H_{012}\approx 0$) on our theory, we have to be careful in expanding the r.h.s. of the above charge $Q_{G}$ so that these constraints are not strongly set equal to zero. In other words, we have the following explicit expression:
@@
Using the following non-zero equal-time commutatorsAt the classical level, we should use the Poisson-brackets instead of the canonical commutators. However, since we shall be discussing the BRST-quantized version of our present classical $3D$ theory, we shall stick with the canonical commutators which will be useful for our discussions at the quantum level.
@@
The gauge-fixing terms in the above owe their origin to the co-exterior derivative $\delta=\pm\,*\,d\,*$ of the differential geometry [14-18]. For instance, it can be readily checked that: $\delta\,A^{(1)}=\pm\,*\,d\,*\big{(}A_{\mu}\,dx^{\mu}\big{)}\equiv\pm\big{(}\partial\cdot A\big{)}$ and $\delta\,B^{(2)}=\pm\,*\,d\,*\big{[}\frac{1}{2!}\,B_{\mu\nu}\,\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}\big{]}\equiv\mp\big{(}\partial^{\nu}B_{\nu\mu}\big{)}dx^{\mu}$ where * is the Hodge duality operator on our chosen $3D$ Minkowskian spacetime manifold. Under the infinitesimal local gauge symmetry transformations (2), we obtain the following:

### v4

# A $3D$ Field-Theoretic Model: Discrete Duality Symmetry

@@
Abstract: We demonstrate the discrete duality symmetry between the Abelian 1-form and 2-form basic gauge fields in the context of a three $(2+1)$-dimensional ($3D$) combined system of the field-theoretic model of the free Abelian 1-from and 2-form gauge theories within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism. The classical gauge-fixed Lagrangian density of this theory is generalized to its quantum counterpart as the BRST and co-BRST invariant Lagrangian density. We show clearly the existence of the off-shell nilpotent (co-)BRST symmetry transformations and establish their intimate connection through a set of underlying discrete duality symmetry transformations in our $3D$ BRST-quantized theory. We provide the mathematical basis for the existence of the discrete duality symmetry transformations in our theory through the Hodge duality operator (that is defined on the $3D$ flat Minkowskian spacetime manifold). We briefly mention a bosonic symmetry transformation which is constructed from the anticommutator of the above off-shell nilpotent (co-)BRST symmetry transformations. We lay emphasis on the algebraic structures of the existing continuous and discrete duality symmetry transformations for our $3D$ BRST-quantized theory (where they are treated as operators). We also comment on the appearance of a pseudo-scalar field (with negative kinetic term). This field happens to be one of the possible candidates for the phantom field of the cosmological models.
@@
Keywords: A $3D$ field-theoretic model; combined system of the Abelian 1-form and 2-form gauge theories; first-class constraints; (dual-)gauge symmetries; (co-)BRST symmetries; a bosonic symmetry; discrete duality symmetry; phantom field
@@ 1 Introduction
The study of the higher $p$-form ($p=2,3,4,$...) gauge theories***The most successful theory of THEP is the standard model of particle physics (SMPP) which is based on the interacting non-Abelian 1-form (i.e. $p=1$) gauge theory where there is a stunning degree of agreements between this theory and experiments. However, the conclusive experimental evidence of the masses of the neutrinos has compelled the modern-day practitioners of THEP to go beyond the domain of the validity of SMPP. The ideas behind the (super)string theories are very promising in this direction. has become important and essential due to the appearance of their basic fields in the quantum excitations of (super)strings which are the forefront areas of research activities in the realm of theoretical high energy physics (THEP). The modern developments in the domain of (super)string theories (see, e.g. [1, 2, 3, 4, 5] and references therein) have brought together, in an unprecedented manner, the active researchers in THEP and pure mathematics on a single intellectual platform where both kinds of researchers have benefited from the knowledge and creativity of each-other. In other words, there has been a notable convergence of ideas from the research activities in the realms of THEP and pure mathematics. Such kinds of confluence of ideas have been observed in our studies of the massless and Stückelberg-modified massive Abelian $p$-form (i.e. $p=1,2,3$) gauge theories (see, e.g. [6, 7, 8, 9] and references therein) within the framework of Becchi-Rouet-Stora-Tyutin (BRST) formalism [10, 11, 12, 13] where it has been established that, in the $D=2p$ (i.e. $D=2,4,6$) dimensions of spacetime, these theories turn out to be the field-theoretic examples for Hodge theory. The discrete and continuous symmetry transformations (and corresponding conserved charges) of these theories have been able to provide the physical realizations of the de Rham cohomological operators of differential geometry (see, e.g. [14, 15, 16, 17, 18]) at the algebraic level. In other words, in our research activities, it has been established that there is a convergence of ideas from the physical aspects of the BRST formalism and the mathematical ingredients related to the cohomological operators of differential geometry. It is crystal clear, from our present discussions, that our studies [6, 7, 8, 9] on the above BRST-quantized Abelian theories have been confined only to the even dimensions (i.e. $D=2,4,6$) of the flat Minkowskian spacetime.
@@
The existence of the discrete duality symmetry transformations in our $3D$ theory is very important as they provide (i) a beautiful algebraic relationship between the off-shell nilpotent (co-)BRST symmetry transformations (cf. Sec. 4), and (ii) a direct connection between the fermionic (i.e. nilpotent) (co-)BRST symmetry transformations of our $3D$ theory (cf. A). In our recent couple of papers [19, 20], we have not been able to obtain such kinds of relationships due to the fact that we were unaware of the existence of such beautiful types of discrete duality symmetry transformations. As a consequence, we could not provide the physical realization of the Hodge duality operator (in our earlier works [19, 20]) which plays a key role in the algebraic structures that are satisfied by the de Rham cohomological operators of differential geometry. In our present endeavor, we have given a great deal of importance to the discrete duality symmetry transformations because they provide (i) a deep connection between the Abelian 1-form and 2-form basic gauge fieldsWe would like to lay emphasis on the physical importance of this duality between the basic Abelian 1-form and 2-form gauge fields because this observation [cf. Eqs. (16),(17)] is at the level of the gauge-fixed Lagrangian density [cf. Eq. (12)] which is completely different from the usual Hodge duality between the 1-form and 2-from that exist mathematically on a given $3D$ compact spacetime manifold. [cf. Eq. (3)], and (ii) the physical realization of the Hodge duality * operator [14, 15, 16, 17, 18] in the relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives [i.e. $(\delta)d$].
@@
One of the key upshots of our present endeavor is the observation that there is existence of the massless (i.e. $\Box\widetilde{\phi}=0,\;\Box\phi=0$) (pseudo-)scalar fields [i.e. $(\widetilde{\phi})\phi$] in our theory which carry the kinetic terms that are endowed with the opposite signs. However, despite this obvious distinct difference, they obey the massless Klein-Gordon equations of motion: $\Box\widetilde{\phi}=0,\;\Box\phi=0$. It is interesting to point that the fields with negative kinetic terms have become quite popular in the realm of the cosmological models of the Universe. In our theory, on the symmetry grounds alone, we observe that the pseudo-scalar field ($\widetilde{\phi}$) carries the negative kinetic term. Such kinds of fields are a set of possible candidates for (i) the “phantom” and/or “ghost” fields in the realm of the cyclic, bouncing and self-accelerated cosmological models of the Universe [21, 22, 23, 24, 25], and (ii) the dark energy (see, e.g. [26, 27] and references therein) because these fields automatically lead to the creation of the negative pressure in the theory (which happens to be one of the characteristic features of dark energy). On the contrary, the pure scalar field ($\phi$), with the positive kinetic term, is a normal field which does not lead to the existence of the negative pressure in the theory. Hence, the existence of the pseudo-scalar field (in our theory) is very interesting in the current scenario of research activities that are related to (i) the realms of the cosmological models of the Universe, and (ii) the search for the possible dark energy candidate(s).
@@
In our A, we establish a direct connection between the off-shell nilpotent BRST and co-BRST symmetry transformations by exploiting the theoretical beauty and strength of the discrete duality symmetry transformations of our $3D$ theory.
@@ 2 Preliminaries: Constraint Analysis, (Dual-)Gauge and Discrete Duality Symmetry Transformations
We begin with the $3D$ field-theoretic model for the combined system of the free Abelian 1-form (i.e. $A^{(1)}=A_{\mu}\,dx^{\mu}$) and 2-form [i.e. $B^{(2)}=\frac{1}{2!}\,B_{\mu\nu}(dx^{\mu}\wedge dx^{\nu})$] gauge theories which is described by the following standard Lagrangian densityOur three $(2+1)$-dimensional ($3D$) Minkowskian spacetime is characterized by the flat metric tensor $\eta_{\mu\nu}=\text{diag}\,(+1,-1,-1)$ so that the dot product between the two non-null Lorentz vectors $U_{\mu}$ and $V_{\mu}$ is defined as: $U\cdot V=\eta_{\mu\nu}\,U^{\mu}\,V^{\nu}\equiv U_{0}\,V_{0}-U_{i}\,V_{i}$ where the Greek indices $\mu,\nu,\sigma,...=0,1,2$ stand for the time and space directions and the Latin indices i,j,k,=1,2i,j,k,...=1,2 denote the space directions only. The $3D$ Levi-Civita tensor $\varepsilon_{\mu\nu\sigma}$ is chosen such that $\varepsilon_{012}=\varepsilon^{012}=+1$ and the standard relationships: $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\sigma}=+3!$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\nu\eta}=+2!\,\delta^{\sigma}_{\eta}$, $\varepsilon^{\mu\nu\sigma}\,\varepsilon_{\mu\eta\kappa}=+1!\,(\delta^{\nu}_{\eta}\,\delta^{\sigma}_{\kappa}-\delta^{\nu}_{\kappa}\,\delta^{\sigma}_{\eta})$ are satisfied. We also adopt the convention of the derivative w.r.t. the antisymmetric tensor field as: $\big{(}\partial B_{\mu\nu}/\partial B_{\sigma\kappa}\big{)}=\frac{1}{2!}\,(\delta^{\sigma}_{\mu}\,\delta^{\kappa}_{\nu}-\delta^{\sigma}_{\nu}\,\delta^{\kappa}_{\mu})$, etc.
@@
where the field-strength tensors: $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $H_{\mu\nu\sigma}=\partial_{\mu}B_{\nu\sigma}+\partial_{\nu}B_{\sigma\mu}+\partial_{\sigma}B_{\mu\nu}$ are derived from the 2-form $F^{(2)}=d\,A^{(1)}=\frac{1}{2!}\,F_{\mu\nu}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}$ and 3-form $H^{(3)}=d\,B^{(2)}=\frac{1}{3!}\,H_{\mu\nu\sigma}\,\big{(}dx^{\mu}\wedge dx^{\nu}\wedge dx^{\sigma}\big{)}$, respectively. Here the symbol $d=\partial_{\mu}\,dx^{\mu}$ [with $d^{2}=\frac{1}{2!}\,\big{(}\partial_{\mu}\,\partial_{\nu}-\partial_{\nu}\,\partial_{\mu}\big{)}\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}=0$] denotes the exterior derivative of differential geometry (see, e.g. [14, 15, 16, 17, 18]) and the above 1-form (i.e. $A^{(1)}$) and 2-form (i.e. $B^{(2)}$) define the Lorentz vector gauge field $A_{\mu}$ and the second-rank antisymmetric (i.e. $B_{\mu\nu}=-\,B_{\nu\mu}$) tensor gauge field $B_{\mu\nu}$, respectively. It is the specific feature of the three $(2+1)$-dimensional $(3D)$ Minkowskian spacetime that the kinetic terms for the Abelian 1-form gauge field $A_{\mu}$ and the antisymmetric tensor gauge field $B_{\mu\nu}$ can be expressed in terms of the $3D$ Levi-Civita tensor and the derivative on them as we have expressed these terms in our equation (1). One of the key reasons behind this observation is the fact that, in the $3D$ Minkowskian spacetime, we have: $\frac{1}{12}\,H^{\mu\nu\sigma}\,H_{\mu\nu\sigma}=\frac{1}{2}\,H^{012}\,H_{012}\equiv\frac{1}{2}\,(H_{012})^{2}$. The component $H_{012}$, however, can be written in its covariant form as: $H_{012}=\frac{1}{2}\,\varepsilon^{\mu\nu\sigma}\,\partial_{\mu}B_{\nu\sigma}$ which has been taken into account in our equation (1) to express the kinetic term for the Abelian 2-form gauge field. Similarly, it is straightforward to check that the $3D$ kinetic term for the Abelian 1-form field is: $\frac{1}{4}\,F^{\mu\nu}\,F_{\mu\nu}=-\frac{1}{2}\,\big{(}\varepsilon^{\mu\nu\sigma}\partial_{\nu}A_{\sigma}\big{)}^{2}$. For our combined system of the $3D$ free Abelian 1-form and 2-form gauge theories, the infinitesimal, local and continuous gauge symmetry transformations $(\delta_{g})$ are well-known. These transformations for our $3D$ system are
@@
To determine the first-class constraints, the first-step is to derive the canonical conjugate momenta (i.e. $\Pi^{\mu\nu}_{(B)},\;\Pi^{\mu}_{(A)}$) w.r.t. the basic gauge fields $B_{\mu\nu}$ and $A_{\mu}$, respectively. For the $3D$ version of the Lagrangian density (1), these momenta are:
@@
The above equation demonstrates that we have two sets of primary constraints (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(A)}\approx 0$) on our theory where the symbol $\approx 0$ stands for weakly zero in Dirac’s notation (see, e.g. [28, 29] for details). The requirements of the time-evolution invariance of these primary constraints lead to the determination of the secondary constraints. The following Euler-Lagrange (EL) equations of motion (EoMs)
@@
where $\Pi^{ij}_{(B)}=\frac{1}{2}\,\varepsilon^{0ij}\,H_{012}$ and $\Pi^{i}_{(A)}=-F^{0i}$ are the space components of the conjugate momenta that have been defined in our equation (2). It is obvious that we have four sets of constraints on our theory out of which two sets (i.e. $\Pi^{0i}_{(B)}\approx 0,\;\Pi^{0}_{(A)}\approx 0$) are the primary constraints and the rest of the two (i.e. $\partial_{j}\Pi^{ji}_{(B)}\approx 0,\;\partial_{i}\Pi^{i}_{(A)}\approx 0$) are the secondary constraints [28, 29]. Since all these constraints are expressed in terms of the components of the conjugate momenta (2), all of them commute among themselves. Hence, they belong to the first-class variety according to Dirac’s prescription for the classification scheme of constraints [28, 29]. These first-class constraints generate the infinitesimal gauge symmetry transformations (2) which we explain in the following paragraph.
According to Noether’s theorem, the existence of the continuous symmetry transformations always leads to the derivations of the Noether conserved currents and corresponding charges. Due to the infinitesimal gauge symmetry transformations in our equation (2), we find the following expression for the Noether gauge current $J^{\mu}_{(G)}$ for our $3D$ field-theoretic combined system of the free Abelian 1-form and 2-form theory
@@
Since we have constraints (i.e. $F^{00}\approx 0,\;\frac{1}{2}\,\varepsilon^{00i}\,H_{012}\approx 0$) on our theory, we have to be careful in expressing the r.h.s. of the above charge $Q_{G}$ so that these constraints are not strongly set equal to zero. In other words, we have the following explicit expression:
@@
Using the following non-zero equal-time commutators§§§At the classical level, we should use the Poisson-brackets instead of the canonical commutators. However, since we shall be discussing the BRST-quantized version of our present classical $3D$ theory, we shall stick with the canonical commutators which will be useful for our discussions at the quantum level.
@@
The gauge-fixing terms in the above owe their origin to the co-exterior derivative $\delta=\pm\,*\,d\,*$ of the differential geometry [14, 15, 16, 17, 18]. For instance, it can be readily checked that: $\delta\,A^{(1)}=\pm\,*\,d\,*\big{(}A_{\mu}\,dx^{\mu}\big{)}\equiv\pm\big{(}\partial\cdot A\big{)}$ and $\delta\,B^{(2)}=\pm\,*\,d\,*\big{[}\frac{1}{2!}\,B_{\mu\nu}\,\big{(}dx^{\mu}\wedge dx^{\nu}\big{)}\big{]}\equiv\mp\big{(}\partial^{\nu}B_{\nu\mu}\big{)}dx^{\mu}$ where * is the Hodge duality operator on our chosen $3D$ Minkowskian spacetime manifold. Under the infinitesimal local gauge symmetry transformations (2), we obtain the following: